### v1

@@ Abstract
3D urban scene reconstruction and modelling is a crucial research area in remote sensing with numerous applications in academia, commerce, industry, and administration. Recent advancements in view synthesis models have facilitated photorealistic 3D reconstruction solely from 2D images. Leveraging Google Earth imagery, we construct a 3D Gaussian Splatting model of the Waterloo region centered on the University of Waterloo and are able to achieve view-synthesis results far exceeding previous 3D view-synthesis results based on neural radiance fields which we demonstrate in our benchmark. Additionally, we retrieved the 3D geometry of the scene using the 3D point cloud extracted from the 3D Gaussian Splatting model which we benchmarked against our Multi-View-Stereo dense reconstruction of the scene, thereby reconstructing both the 3D geometry and photorealistic lighting of the large-scale urban scene through 3D Gaussian Splatting.
@@ 3.5.1 Rasterization
The tile-based rasterizer tiles the image into $16\times 16$. For each tile, a view frustrum is projected into the 3D scene. 3D Gaussians are accumulated/assigned per-tile according to overlap to the view frustrum, and are projected into 2D and via projection of its mean $\mu$ and covariance matrix $\Sigma$. These are given by
$\mu^{\prime}=k(W\mu)$ (10)
@@
where $W$ is the view transformation and $k$ is the projective transformation. The transformation is a matrix multiplication in the case of a linear camera model such as the pinhole model used with the standard 3DGS model, which results in $\mu^{\prime}=KW\mu$. $J$ is the Jacobian of the affine approximation of the projective transformation ($J=K$ in linear camera models). The third row and third column of $\Sigma^{\prime}$ are then skipped to obtain the 2D covariance matrix in the imaging plane.

### v2

@@ Abstract
Although large-scale remote sensing image based 3D urban scene reconstruction and modelling is crucial for many key applications such as digital twins and smart cities, it is a difficult task due to the uncertainties in heterogeneous datasets and the geometry models. This paper presents a Gaussian splatting based approach for 3D urban scene modeling and geometry retrieval, with the following contributions. First, we develop and implement a 3D Gaussian splatting (3DGS) approach for large-scale 3D urban scene modeling from heterogeneous remote sensing images. Second, we design point cloud densification approach in the proposed 3DGS model to improve the quality of 3D geometry extraction of urban scenes. Leveraging Google Earth imagery of different sensors, the proposed approach is tested on the region of University of Waterloo, demonstrating that the proposed approach greatly improves reconstructed point clouds quality over the other Multi-View-Stereo approaches. Third, we design and conduct extensive experiments on multi-source large-scale Google Earth remote sensing images across ten cities to compare the 3DGS approach with neural radiance field (NeRF) approaches, demonstrating improved view-synthesis results that greatly outperform previous state-of-the-art 3D view-synthesis approaches.
@@ 3.5.1 Rasterization
The tile-based rasterizer tiles the image into $16\times 16$. For each tile, a view frustrum is projected into the 3D scene. 3D Gaussians are accumulated/assigned per-tile according to overlap to the view frustrum, and are projected into 2D and via projection of its covariance matrix $\Sigma$. Starting in homogenous coordinates, this is given by
@@
where $W$ is the view transformation and $J$ is affine approximation of the projective transformation. The projective transformation is a matrix multiplication in homogenous coordinates in the case of a linear camera model such as the pinhole model used with the standard 3DGS model. In which case, $J$ is simply obtained from the intrinsic camera matrix. The third column of $\mu^{\prime}$, and the third row and third column of $\Sigma^{\prime}$ are then skipped to obtain the 2D mean and the 2D covariance matrix in the imaging plane in Cartesian coordinates.